In examining the determinants of income, data were collected regarding the characteristics of 45 adults, and the regression Y = β1 + β2S + β3D + β3 (S*D )+ u was used, where Yis the annual income (in thousands of dollars), S is his/her years of education, and D is a dummy variable = 1 if the adult is female.


If you get = 26.3 + 2.94S - 0.76D+ 0.34(SD) when you run the regression, which of the following is true?

  • A.

    The regression for males is   = 26.3 + 2.94S - 0.76D+ 0.34(SD)

  • B.

    The regression for females is   = 25.54 + 3.28S

  • C.

    The regression for males is   = 25.54 + 2.94S

  • D.

    None of these regressions are correct.
Answer #1

Answer: (B) The regression for females is Y = 25.54 + 3.28S

The regression equation is:

Y = 26.3 + 2.94S - 0.76D + 0.34(S ∙ D)

>> The regression equation for male is:

Y = 26.3 + 2.94S - 0.76*0 + 0.34(S*0)

Y = 26.3 + 2.94S

>> The regression equation for female is:

Y = 26.3 + 2.94S - 0.76*1 + 0.34(S*1)

Y = 25.54 + 3.28S